A sample of hydrogen gas (H2) has a volume of 17.4 L at a temperature of 0°C and a pressure of 1.60 atm. The number of moles of H2 molecules of hydrogen in the sample is
1
Expert's answer
2013-02-28T11:18:50-0500
Clapeyron-Mendeleev law equation can be used foranswering this question:

PV = nRT

n is amount = ?
P is pressure = 1.60 atm = 162.12 kPa
V is volume = 17.4 L
R is gas constant = 8.31
T is temperature = 0°C = 273K

n = PV/RT

n = 162.12 * 17.4 / (8.31*273) = 1.24 mol